# Bounding User Contributions for User-Level Differentially Private Mean Estimation

### v4

@@ I Introduction
In this article, we work on the fundamental problem of processing bounded, potentially vector-valued samples in a dataset, for the release of a private estimate of the sample mean. In particular, we work within the framework of “user-level” differential privacy [userlevel, hetero], which is a generalization of the now widely adopted framework of differential privacy (DP) [dworkroth, vadhan2017] for the design and analysis of privacy-preserving algorithms. Loosely, user-level DP guarantees the privacy of a “user”, who could contribute more than one sample, by ensuring the statistical indistinguishability of outputs of the algorithm to changes in the user’s samples. User-level DP has practical relevance for inference tasks on real-world datasets, such as traffic datasets, datasets of user expenditures, and time series data, where different users contribute potentially different numbers of samples (data heterogeneity) [usereg3, usereg4], and more recently, in federated learning (FL) applications (see, e.g., [kairouz-survey, Sec. 4]).
@@
In this article, we confine our attention to user-level (pure) $\varepsilon$-DP algorithms for mean estimation. A key subroutine in such mechanisms [userlevel, hetero-user-level, hist-user-level, amin] is the preprocessing of the data samples for the release of an estimate of the sample mean, which requires the addition of less noise for privacy, as against releasing a noised version of the true sample mean. Such a preprocessing procedure (or “bounding” or “clipping” strategy [amin]) either drops selected samples, or projects the samples to a “high-probability subset”. While it is usually easy to establish that such mechanisms are differentially private, an analysis of their “utility”, or the error in estimation of the true statistic, often relies on distributional assumptions about the dataset.
In this work, following [dp_spcom, dp_preprint, tit-preprint], we define and explicitly compute the worst-case error, over all datasets, of general preprocessing (or bounding) strategies. This error metric is natural in settings with arbitrarily correlated data, where a user potentially ascribes his/her error tolerance to the worst dataset that the statistic is computed on. Furthermore, this error metric is distribution-independent and, as we show, is computable under data heterogeneity too. We then explicitly identify the bounding strategy that results in the smallest worst-case error. Interestingly, we also observe from experimental studies that for scalar samples, our clipping strategy also gives rise to much smaller errors on average compared to that in [amin], for selected dataset sizes, when the samples are drawn i.i.d. according to common distributions.
@@ II-B Problem Formulation
Let $L$ be the number of users present in the dataset. For every user $\ell\in[L]$, let the number of records contributed by the user be denoted by $m_{\ell}$, and set $m^{\star}:=\max_{\ell\in[L]}m_{\ell}$ and $m_{\star}:=\min_{\ell\in[L]}m_{\ell}$. We assume that $L$ and the collection $\{m_{\ell}:\ell\in[L]\}$ are known. Now, for a given user $\ell\in[L]$, let $\left\{\mathbf{x}_{j}^{(\ell)}:j\in[m_{\ell}]\right\}$ denote the collection of (potentially arbitrary) bounded samples contributed by the user, where each $\mathbf{x}_{j}^{(\ell)}\in\mathbb{R}^{d}$, for some dimension $d\geq 1$. We assume, as is common for most applications [dpsgd, dp_spcom], that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}$, for all $\ell\in[L]$, $j\in[m_{\ell}]$, where $U>0$ is known. Call the dataset as $\mathcal{D}=\left\{\left(\ell,\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$.
@@ II-C Differential Privacy
We use the terms “sensitivity” and “user-level sensitivity” interchangeably. The next result is well-known [dwork06, Prop. 1].
@@ II-D The Worst-Case Error Metric
All through, in this paper, we shall work with user-level $\varepsilon$-DP mechanisms that add a suitable amount of Laplace noise that is tailored to the sensitivity of the function used as an estimator of the sample mean $f$. Consider a mechanism $M$ for the user-level $\varepsilon$-DP release of the statistic $f$. The canonical structure of $M$ (see [staircase, Footnote 1], [opt]) is: $M(\mathcal{D})=\overline{f}(\mathcal{D})+{\mathbf{Z}},$ for some estimator $\overline{f}$ of $f$, with user-level sensitivity $\Delta_{\overline{f}}$, with $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$, with $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}\left(\Delta_{\overline{f}}/\varepsilon\right)$, $i\in[d]$.
@@
Clearly, the expression above is an upper bound on the $\ell_{1}$-error $E^{(1)}_{M}:=\max_{\mathcal{D}\in\mathsf{D}}\mathbb{E}[|f(\mathcal{D})-M(\mathcal{D})|]$, via the triangle inequality. When the dimension $d$ is clear from the context, we simply denote $E_{\overline{f}}^{(d)}$ as $E_{\overline{f}}$. The distribution-independent expression in (1) conveniently captures the errors due to bias (the first term) and due to noise addition for privacy (the second term); a similar such error measure that separates the bias and noise errors was employed in [amin].
@@ III-A On Clipping Strategies
Note that the class $\mathsf{B}$ of estimators $\overline{f}$ as above captures those estimators obtained by dropping selected samples $\mathbf{x}_{j}^{(\ell)}$ (by setting $a_{j}^{(\ell)}=b_{j}^{(\ell)}=0$ for those samples) and those obtained by projecting samples onto an $\ell_{1}$-bounded subset of $\Delta_{U}$, in addition to strategies that perform a combination of dropping and projection. This class of estimators hence includes several common estimators of the sample mean used in works such as [userlevel, dp_preprint, tit-preprint].
@@ III-B On Worst-Case Errors of Clipping Strategies
We have that $f^{\star}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizes $E_{\overline{f}}$, where $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{Um_{\ell}+T_{\varepsilon}}{2m_{\ell}},U\right\},$ if $d=1$, and $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{T_{\varepsilon}}{m_{\ell}},U\right\}$, for any $d\geq 2$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Some remarks are in order. First, note that the optimal bounding strategy $f^{\star}$ clips sample values based only on the number of contributions of each user $\ell\in[L]$. Furthermore, note that the interval of projection is determined by $T_{\varepsilon}$, which is quite similar in structure to the optimal clipping threshold $T$ in [amin] for item-level DP, which is the (privately estimated) $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample value.
@@
For any $d\geq 1$, there exists an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$, which obeys $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
@@
Consider any optimal estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$, and suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}>U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$]. The proof when we suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}<U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$], is similar, and is omitted. Let $\delta=(a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})})-U$. By setting $b_{j}^{(\tilde{\ell})}\leftarrow b_{j}^{(\tilde{\ell})}-\delta$, we observe that $\beta^{(d)}(\overline{f})$ remains unchanged, while $\eta^{(d)}(\overline{f})$ either remains unchanged or strictly decreases by $\delta>0$. ∎
Hence, to obtain the explicit structure of an estimator $\overline{f}$ that minimizes $E_{\overline{f}}$, it suffices to focus estimators $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ with $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$, $j\in[m_{\ell}]$. For $\ell\in[L]$, let $S^{(\ell)}:=\sum_{j\leq m_{\ell}}a_{j}^{(\ell)}$, for any given estimator $\overline{f}$. Then, for $d=1$, there exists an estimator $f^{\star}$ minimizing $E^{(1)}_{\overline{f}}$ that satisfies
$E^{(1)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{d}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)\right);$ (4)
likewise, for any $d\geq 2$, there exists an estimator $f^{\star}$ minimizing $E^{(d)}_{\overline{f}}$ that satisfies
$E^{(d)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{2d}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-S^{(\ell)}\right)\right).$ (5)
We now prove Theorem III.1.
@@
We prove the theorem in detail for the case when $d=1$; the setting of $d\geq 2$ is quite similar, and is omitted. We begin with the expression in (4); note that our task now is simply to identify the optimal parameters $\{S^{(\ell)}:\ \ell\in[L]\}$ of $f^{\star}$ in (4). Once these parameters are derived, we simply set $a_{j}^{(\ell)}:=S^{(\ell)}/m_{\ell}=U-b_{j}^{(\ell)}$ (via Lemma III.4), for each $\ell\in[L]$ and $j\in[m_{\ell}]$. In the expression in (4), let us set $\tau:=\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)$. We hence need to solve the following constrained optimization problem:
${\text{minimize}}\quad h(\{S^{(\ell)}\}):=\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{d\tau}{\varepsilon}\right)$
@@
The optimization problem in (6) is a linear programming problem. By standard arguments via the necessity of the KKT conditions [boyd, Sec. 5.5.3], there must exist reals $\lambda_{\tau}\geq 0$ and $\lambda_{\ell},\ \mu_{\ell}\geq 0$, for each $\ell\in[L]$ (or Lagrange multipliers), such that the function
$\mathcal{L}(\{S^{(\ell)}\},\$$\lambda_{S},\{\lambda_{\ell},\mu_{\ell}\}):=\sum_{\ell\leq L}S^{(\ell)}+\frac{S}{\varepsilon}-\lambda_{S}S\ +$
@@
We claim that the assignment $\tau^{\star}=T_{\varepsilon}$ and $S^{(\ell),\star}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2},0\right\}$ satisfies the conditions above, for an appropriate choice of $\lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}$ values. Indeed, for $\ell\in[L]$, if $S^{(\ell),\star}=0$, we set $\lambda_{\ell}=0$ and $\mu_{\ell}=1$; else, we set $\lambda_{\ell}=\frac{1}{2}$ and $\mu_{\ell}=0$. ∎
@@
and the the optimal worst-case error for any $d\geq 2$ is
$E^{\text{OPT, $(d)$}}(\varepsilon)=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{{Um_{\ell}-T_{\varepsilon}},0\right\}+\frac{2dT_{\varepsilon}}{\varepsilon}\right).$ (8)
We end this section with a couple of remarks. First, observe from (7) and (8) that, in the limit as $\varepsilon\to 0$, the optimal worst-case error equals the worst-case bias error, which in turn equals $U/2$, when $d=1$, and $U$, when $d\geq 2$.

### v5

@@ I Introduction
In this article, we work on the fundamental problem of processing bounded, potentially vector-valued samples in a dataset, for the release of a private estimate of the sample mean. In particular, we work within the framework of “user-level” differential privacy [15, 6], which is a generalization of the now widely adopted framework of differential privacy (DP) [9, 19] for the design and analysis of privacy-preserving algorithms. Loosely, user-level DP guarantees the privacy of a “user”, who could contribute more than one sample, by ensuring the statistical indistinguishability of outputs of the algorithm to changes in the user’s samples. User-level DP has practical relevance for inference tasks on real-world datasets, such as traffic datasets, datasets of user expenditures, and time series data, where different users contribute potentially different numbers of samples (data heterogeneity) [20, 17], and more recently, in federated learning (FL) applications (see, e.g., [14, Sec. 4]).
@@
In this article, we confine our attention to user-level (pure) $\varepsilon$-DP algorithms for mean estimation. A key subroutine in such mechanisms [15, 7, 16, 2] is the preprocessing of the data samples for the release of an estimate of the sample mean, which requires the addition of less noise for privacy, as against releasing a noised version of the true sample mean. Such a preprocessing procedure (or “bounding” or “clipping” strategy [2]) either drops selected samples, or projects the samples to a “high-probability subset”. While it is usually easy to establish that such mechanisms are differentially private, an analysis of their “utility”, or the error in estimation of the true statistic, often relies on distributional assumptions about the dataset.
In this work, following [13, 3, 18], we define and explicitly compute the worst-case error, over all datasets, of general preprocessing (or bounding) strategies. This error metric is natural in settings with arbitrarily correlated data, where a user potentially ascribes his/her error tolerance to the worst dataset that the statistic is computed on. Furthermore, this error metric is distribution-independent and, as we show, is computable under data heterogeneity too. We then explicitly identify the bounding strategy that results in the smallest worst-case error. Interestingly, we also observe from experimental studies that for scalar samples, our clipping strategy also gives rise to much smaller errors on average compared to that in [2], for selected dataset sizes, when the samples are drawn i.i.d. according to common distributions.
@@ II-B Problem Formulation
Let $L$ be the number of users present in the dataset. For every user $\ell\in[L]$, let the number of records contributed by the user be denoted by $m_{\ell}$, and set $m^{\star}:=\max_{\ell\in[L]}m_{\ell}$ and $m_{\star}:=\min_{\ell\in[L]}m_{\ell}$. We assume that $L$ and the collection $\{m_{\ell}:\ell\in[L]\}$ are known. Now, for a given user $\ell\in[L]$, let $\left\{\mathbf{x}_{j}^{(\ell)}:j\in[m_{\ell}]\right\}$ denote the collection of (potentially arbitrary) bounded samples contributed by the user, where each $\mathbf{x}_{j}^{(\ell)}\in\mathbb{R}^{d}$, for some dimension $d\geq 1$. We assume, as is common for most applications [1, 13], that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}$, for all $\ell\in[L]$, $j\in[m_{\ell}]$, where $U>0$ is known. Call the dataset as $\mathcal{D}=\left\{\left(\ell,\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$.
@@ II-C Differential Privacy
We use the terms “sensitivity” and “user-level sensitivity” interchangeably. The next result is well-known [8, Prop. 1].
@@ II-D The Worst-Case Error Metric
All through, in this paper, we shall work with user-level $\varepsilon$-DP mechanisms that add a suitable amount of Laplace noise that is tailored to the sensitivity of the function used as an estimator of the sample mean $f$. Consider a mechanism $M$ for the user-level $\varepsilon$-DP release of the statistic $f$. The canonical structure of $M$ (see [11, Footnote 1], [12]) is: $M(\mathcal{D})=\overline{f}(\mathcal{D})+{\mathbf{Z}},$ for some estimator $\overline{f}$ of $f$, with user-level sensitivity $\Delta_{\overline{f}}$, with $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$, with $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}\left(\Delta_{\overline{f}}/\varepsilon\right)$, $i\in[d]$.
@@
Clearly, the expression above is an upper bound on the $\ell_{1}$-error $E^{(1)}_{M}:=\max_{\mathcal{D}\in\mathsf{D}}\mathbb{E}[|f(\mathcal{D})-M(\mathcal{D})|]$, via the triangle inequality. When the dimension $d$ is clear from the context, we simply denote $E_{\overline{f}}^{(d)}$ as $E_{\overline{f}}$. The distribution-independent expression in (1) conveniently captures the errors due to bias (the first term) and due to noise addition for privacy (the second term); a similar such error measure that separates the bias and noise errors was employed in [2].
@@ III-A On Clipping Strategies
Note that the class $\mathsf{B}$ of estimators $\overline{f}$ as above captures those estimators obtained by dropping selected samples $\mathbf{x}_{j}^{(\ell)}$ (by setting $a_{j}^{(\ell)}=b_{j}^{(\ell)}=0$ for those samples) and those obtained by projecting samples onto an $\ell_{1}$-bounded subset of $\Delta_{U}$, in addition to strategies that perform a combination of dropping and projection. This class of estimators hence includes several common estimators of the sample mean used in works such as [15, 3, 18].
@@ III-B On Worst-Case Errors of Clipping Strategies
We have that $f^{\star}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizes $E_{\overline{f}}$, where $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{Um_{\ell}+T_{\varepsilon}}{2m_{\ell}},U\right\},$ if $d=1$, and $a_{j}^{(\ell)}=0\ \text{ and }b_{j}^{(\ell)}=\min\left\{\frac{T_{\varepsilon}}{m_{\ell}},U\right\}$, for any $d\geq 2$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Some remarks are in order. First, note that the optimal bounding strategy $f^{\star}$ clips sample values based only on the number of contributions of each user $\ell\in[L]$. Furthermore, note that the interval of projection is determined by $T_{\varepsilon}$, which is quite similar in structure to the optimal clipping threshold $T$ in [2] for item-level DP, which is the (privately estimated) $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample value.
@@
For $d=1$, there exists an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$, which obeys $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$. Furthermore, for $d\geq 2$, an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$ must set $a_{j}^{(\ell)}=0$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
@@
Consider any optimal estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$, and suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}>U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$]. The proof when we suppose that $a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})}<U$, for some $\tilde{\ell}\in[L]$ and $j\in[m_{\tilde{\ell}}$], is similar, and is omitted. Let $\delta=(a_{j}^{(\tilde{\ell})}+b_{j}^{(\tilde{\ell})})-U$. By setting $b_{j}^{(\tilde{\ell})}\leftarrow b_{j}^{(\tilde{\ell})}-\delta$, we observe that $\beta^{(d)}(\overline{f})$ remains unchanged, while $\eta^{(d)}(\overline{f})$ either remains unchanged or strictly decreases by $\delta>0$. Moreover, for $d\geq 2$, note from Proposition III.1 that increasing any $a_{j}^{(\ell)}$ from 0 to a positive value can only strictly increase $E_{\overline{f}}$. ∎
Hence, to obtain the explicit structure of an estimator $\overline{f}$ that minimizes $E_{\overline{f}}$, it suffices to focus estimators $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ with $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$, $j\in[m_{\ell}]$. Then, for $d=1$, there exists an estimator $f^{\star}$ minimizing $E^{(1)}_{\overline{f}}$ that satisfies
$E^{(1)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{1}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)\right),$ (4)
where for $\ell\in[L]$, we let $S^{(\ell)}:=\sum_{j\leq m_{\ell}}a_{j}^{(\ell)}$, for any given estimator $\overline{f}$. Furthermore, for any $d\geq 2$, there exists an estimator $f^{\star}$ minimizing $E^{(d)}_{\overline{f}}$ that satisfies
$E^{(d)}_{f^{\star}}=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\overline{S}^{(\ell)}+\frac{2d}{\varepsilon}\cdot\max_{\ell\leq L}\left(Um_{\ell}-\overline{S}^{(\ell)}\right)\right),$ (5)
where for $\ell\in[L]$, we let $\overline{S}^{(\ell)}:=\sum_{j\leq m_{\ell}}\left(U-b_{j}^{(\ell)}\right)$, for any given estimator $\overline{f}$. We now prove Theorem III.1.
@@
We prove the theorem in detail for the case when $d=1$; the setting of $d\geq 2$ is quite similar and the proof is hence is omitted. We begin with the expression in (4); note that our task now is simply to identify the optimal parameters $\{S^{(\ell)}:\ \ell\in[L]\}$ of $f^{\star}$ in (4). Once these parameters are derived, we simply set $a_{j}^{(\ell)}:=S^{(\ell)}/m_{\ell}=U-b_{j}^{(\ell)}$ (via Lemma III.4), for each $\ell\in[L]$ and $j\in[m_{\ell}]$. In the expression in (4), let us set $\tau:=\max_{\ell\leq L}\left(Um_{\ell}-2S^{(\ell)}\right)$. We hence need to solve the following constrained optimization problem:
${\text{minimize}}\quad h(\{S^{(\ell)}\}):=\left(\sum_{\ell\leq L}S^{(\ell)}+\frac{\tau}{\varepsilon}\right)$
@@
The optimization problem in (6) is a linear programming problem. By standard arguments via the necessity of the KKT conditions [4, Sec. 5.5.3], there must exist reals $\lambda_{\tau}\geq 0$ and $\lambda_{\ell},\ \mu_{\ell}\geq 0$, for each $\ell\in[L]$ (or Lagrange multipliers), such that the function
$\mathcal{L}(\{S^{(\ell)}\},\$$\lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}):=\sum_{\ell\leq L}S^{(\ell)}+\frac{S}{\varepsilon}-\lambda_{\tau}\tau\ +$
@@
We claim that the assignment $\tau^{\star}=T_{\varepsilon}$ and $S^{(\ell),\star}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2},0\right\}$ satisfies the conditions above, for an appropriate choice of $\lambda_{\tau},\{\lambda_{\ell},\mu_{\ell}\}$ values. Indeed, for $\ell\in[L]$, if $S^{(\ell),\star}=0$, we set $\lambda_{\ell}=0$ and $\mu_{\ell}=1$; else, we set $\lambda_{\ell}=\frac{1}{2}$ and $\mu_{\ell}=0$.
@@
and the the optimal worst-case error for any $d\geq 2$ is such that
$E^{\text{OPT, $(d)$}}(\varepsilon)$
$=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\max\left\{{Um_{\ell}-T_{\varepsilon}},0\right\}+\frac{2dT_{\varepsilon}}{\varepsilon}\right).$ (8)
We end this section with a couple of remarks. First, observe from (7) and (8) that, in the limit as $\varepsilon\to 0$, the optimal worst-case error equals the worst-case bias error, which in turn equals $U/2$, for $d=1$, and $U$, when $d\geq 2$.